What is the electric field (E) produced by the source charge (+10 Coul.) at a distance of 0.01 m? (ke = 8.99 X 10^9 N·m^2/C^2)
will rate thank you!
Answer #1

Electric field

E=\frac{1}{4\pi \epsilon _{0}}\frac{q}{r^{2}}

\frac{1}{4\pi \epsilon _{0}}=8.99*10^{9}Nm^{2}/C^{2}   q = 10 C,    r = 0.01 m

E=\frac{8.99*10^{9}*10}{0.01^{2}}=8.99*10^{14}N/C